# Simulation-Based Inference for Direction Reconstruction of Ultra-High-Energy Cosmic Rays with Radio Arrays

### v1

@@ Abstract
Ultra-high-energy cosmic-ray (UHECR) observatories require unbiased direction reconstruction to enable multi-messenger astronomy with sparse, nanosecond-scale radio pulses. Explicit likelihood methods often rely on simplified models, which may bias results and understate uncertainties. We introduce a simulation-based inference pipeline that couples a physics-informed graph neural network (GNN) to a normalizing-flow posterior within the Learning the Universe Implicit Likelihood Inference framework. Each event is seeded by an analytic plane-wavefront fit; the GNN refines this estimate by learning spatiotemporal correlations among antenna signals, and its frozen embedding conditions an eight-block autoregressive flow that returns the full Bayesian posterior. Trained on about $8,000$ realistic UHECR air-shower simulations generated with the ZHAireS code, the posteriors are temperature-calibrated to meet empirical coverage targets. We demonstrate a sub-degree median angular resolution on test UHECR events, and find that the nominal 68% highest-posterior-density contours capture $71\%\pm 2\%$ of true arrival directions, indicating a mildly conservative uncertainty calibration. This approach provides physically interpretable reconstructions, well-calibrated uncertainties, and rapid inference, making it ideally suited for upcoming experiments targeting highly inclined events, such as GRAND, AugerPrime Radio, IceCube-Gen2, RNO-G, and BEACON.
@@ I Introduction
Ultra-high-energy cosmic rays (UHECRs) are the most energetic particles detected in nature [1, 2], but their origins and propagation mechanisms remain poorly understood [3, 4, 5]. Accurate detection and reconstruction of these rare events have significant implications for fundamental physics and astronomy [6, 7]. In particular, precisely reconstructing the arrival direction of UHECRs is essential: it mitigates false detections and provides a foundation for inferring related air-shower properties such as primary particle energy and composition [8, 9, 10]. Additionally, directional accuracy is critical for multi-messenger astrophysics; for instance, distinguishing showers arriving from below the horizon serves as a clear indicator of neutrino-induced events rather than conventional cosmic-ray interactions [11].
@@
Next-generation observatories are rising to this challenge by focusing on radio detection techniques, which offer all-sky coverage and high-duty cycles [12]. Large-scale projects such as the planned Giant Radio Array for Neutrino Detection (GRAND) [13], AugerPrime Radio [14], and BEACON [15], as well as in-ice radio arrays like RNO-G [16] and IceCube-Gen2 [17], are designed to instrument vast areas with antenna networks to capture the nanosecond radio pulses from extensive air showers. These instruments are particularly sensitive to inclined (nearly horizontal) air showers, which produce broad radio footprints on the ground and are a prime channel for detecting Earth-skimming neutrinos and high-energy cosmic rays near the horizon. With their large apertures and exquisite angular resolutions, radio arrays open a new window for UHECR and neutrino detection, provided that we have robust methods to reconstruct the event geometry from the recorded radio signals [18].
@@
The detection of UHECRs using radio arrays is now a well-established technique. Modern radio arrays—such as LOPES [19], CODALEMA [20], LOFAR [21], AERA [22], and TREND [23]; reconstruct the incoming direction of UHECRs by fitting antenna positions and measured trigger times to a geometric model of the shower wavefront.
@@
A reconstruction pipeline has recently been proposed [8] for GRAND that leverages the amplitude distribution of the radio signal in a physics-informed, multi-stage approach. The procedure begins with a plane wavefront (PWF) timing fit to obtain an initial estimate of the arrival direction. This initial estimate is subsequently refined by introducing a mild curvature represented by a spheroidal surface wavefront model. In the final step, the lateral distribution of the radio amplitude is introduced to constrain the shower core position and axis orientation precisely. This hierarchical approach yields robust reconstruction performance, especially for air showers at high zenith angles [24].
@@
Although extremely promising, this method also exposes the limits of current reconstruction pipelines. Its successive modules—planar fitting, curvature correction, and amplitude weighting—stand on model-dependent assumptions that could introduce biases. Traditional explicit likelihood methods, such as formulating a likelihood for observed arrival times given a trial geometry and exploring it via Markov Chain Monte Carlo, are conceptually appealing but practically prohibitive for large arrays and complex wavefront structures [24]. Template-matching schemes confront similar scalability bottlenecks; reproducing the full span of arrival directions, primary energies, and atmospheric states demands vast libraries of simulated air-shower footprints, whose storage and compute requirements grow steeply with each gain in angular or temporal resolution [25].
@@
Equally important, conventional reconstruction algorithms often fail to rigorously quantify uncertainties; while they may yield a best-fit arrival direction, recovering the full posterior distribution remains non-trivial and is frequently overlooked. The recent analytical PWF solution in Ref. [26] addresses part of this gap by providing a closed-form direction estimator with error estimates for the planar approximation. Yet, a planar wavefront model can only capture so much. To fully exploit the rich information in the radio signal pattern, more flexible inference approaches are needed.
@@
A promising route to overcome the limitations of explicit likelihood methods is simulation-based inference (SBI) [27], also known as implicit or likelihood-free inference. Rather than relying on an analytic likelihood function, SBI leverages high-fidelity simulations of the physics process to directly learn the relationship between model parameters (e.g. shower direction) and observables (antenna times and signals). In a Bayesian context, SBI methods aim to directly estimate the posterior distribution of parameters by training machine learning models on simulated data, thereby “learning” an approximate likelihood function or posterior emulator (see Ref. [28, 29] and references therein).
This approach is well-suited to astroparticle problems like UHECR air showers, where we have extremely sophisticated Monte Carlo simulators (e.g., ZHAireS [30], CoREAS [31]) but no simple closed-form likelihood for the radio observables. By training on simulations, SBI can in principle capture the complex, nonlinear dependencies in radio data (including realistic noise and detector effects) without requiring simplifying assumptions in the inference stage [27].
@@
In recent years, SBI techniques have gained traction across astrophysics and cosmology, often outperforming traditional analysis methods [32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Examples range from inferring cosmological parameters from galaxy surveys to decoding gravitational wave signals, where neural density estimators have achieved faster and more accurate results than classical MCMC in high-dimensional parameter spaces (See also Ref. [28] for a detailed discussion). These successes underscore the appeal of SBI; it provides fully Bayesian posteriors with uncertainty quantification for problems that are otherwise intractable with explicit likelihoods [29]. Importantly, likelihood-free approaches can naturally handle complicated selection effects or “cuts” on the data by applying the same cuts to simulated data, a notable advantage when dealing with complex experimental setups.
@@
One such framework is Learning the Universe—Implicit Likelihood Inference (LtU-ILI) [29]. LtU-ILI provides an end-to-end SBI platform that incorporates adaptable neural encoders, inference strategies, and rigorous validation tools. A notable feature of LtU-ILI is its support for complex data representations through custom “embedding” networks—including convolutional networks for image-like data and graph neural networks (GNNs) for relational data.
@@
This is particularly pertinent for radio detector arrays such as GRAND, where the spatial configuration of antennas and their trigger times can be naturally modeled as a graph (with antennas as nodes and edges connecting neighbor detectors). Such a representation enables graph neural networks to effectively capture spatiotemporal correlations in the developing air-shower front. Moreover, embedded in a Bayesian workflow, LtU-ILI audits its posteriors with coverage tests to verify statistical calibration [29]. The result is an inference pipeline that remains computationally tractable while preserving clear physical meaning.
@@ II.1 Physical Modeling of Extensive Air Showers
When an UHECR primary strikes the atmosphere, it initiates a cascade of secondary particles known as an extensive air shower (EAS). The primary collision produces a multitude of daughter particles that branch into a hadronic component (secondary mesons like pions and their decay products) and an electromagnetic component (e.g., photons from $\pi^{0}$ decay feeding $e^{+}e^{-}$ cascades) [12]. This process continues until the primary’s energy is dissipated, yielding a shower of billions of particles spread over kilometers by the time it reaches ground level. Accurately modeling these cascades is essential, as the spatial, temporal, and energy distributions of the secondaries encode the nature of the primary particle [30]. In particular, the arrival direction of the primary can be inferred from the pattern of particle arrival times or atmospheric emissions, providing critical clues to the UHECR’s origin [8, 24, 26].
Our work focuses on radio-detection of air showers, a technique that has matured over the past two decades. As the EAS evolves, its charged component experiences geomagnetic deflection and charge separation in the Earth’s magnetic field, driving transverse currents that emit coherent radio pulses. Additionally, a negative charge excess builds up in the shower front (due to electrons knocking out ambient electrons from air molecules), leading to Askaryan radiation [12, 11]. The total radio signal is a superposition of these geomagnetic and Askaryan contributions. Modern experiments like LOFAR [21], AERA [22], and others have demonstrated that radio arrays can measure UHECR showers with impressive precision, rivaling conventional particle detectors in reconstructing the shower’s arrival direction and even its energy and shower maximum ($X_{\rm max}$). Radio detection thus provides a robust method to observe UHECR events over extensive areas and under virtually all weather conditions. This is achieved by recording the nanosecond-scale electromagnetic pulses generated by air showers using simple dipole antennas [13].
@@ II.2 Synthetic Data Generation and Radio-Signal Formation
To infer the arrival directions of UHECRs, we utilize detailed Monte Carlo simulations of extensive air showers. These simulations [42] were developed by the GRAND collaboration to facilitate the design and optimization of detection algorithms for GRANDproto300 [43], a prototype array comprising $\sim 300$ antennas, presently under construction in China’s Gobi Desert.
@@
The simulations were constructed with the ZHAireS code  [^1] http://aires.fisica.unlp.edu.ar/zhaires/, a simulation package built on the AIRES [45, 46] air-shower engine and extended to calculate radio emission in detail. ZHAireS [30] tracks the full shower development in the atmosphere and computes the radiated electromagnetic fields from charged particle motion based on first principles. Each simulation propagates particles through the atmosphere in small steps, using Liénard-Wiechert potentials to calculate the emitted electric field from each track segment.
@@
This simulator inherently accounts for interference effects among the myriad moving charges and includes realistic propagation conditions, such as the altitude-dependent atmospheric refractive index. As a result, ZHAireS can predict the complete radio pulse waveform at any observer position, in both time and frequency domains. In addition, ZHAireS can simulate the response of a given antenna array to the shower: for each simulated event, it outputs the electric field trace expected at each antenna, from which one can extract observables like signal peak times and electric field (E-field) amplitudes (see, e.g., Ref. [47] and references therein).
@@
We generated antenna-level waveforms using ZHAireS for downward-going air showers initiated by protons and iron nuclei, spanning energies from 0.4 to 4 EeV. Our simulations span zenith angles $37^{\circ}\leq\theta\leq 87^{\circ}$ with uniform coverage in azimuth ($0^{\circ}-360^{\circ}$). All runs adopt a geomagnetic field and atmospheric model (air density/refractivity) appropriate for the GRANDproto300 site in the Gobi desert of China.
To emulate the GRANDproto300 digitizer timing precision, we first pass the clean ZHAireS E-field traces through a filter. Specifically, channels whose peak E-field amplitude lies below $60~\mu\mathrm{V/m}$ are excised, and only events illuminating at least five antennas are retained. For every surviving trace, we then compute the analytic envelope via a Hilbert transform, locate its maximum to define the pulse time stamp, and finally perturb that time stamp with a zero-mean Gaussian deviate of $\sigma=5\mathrm{~ns}$, which corresponds to the GPS-synchronisation jitter estimated for GRANDproto300 [48]. No additional timing errors arising from Galactic foregrounds, receiver noise, or gain calibration drift are introduced at this stage. This ordering cleanly decouples amplitude systematics from timing jitter; we will tackle the former in a forthcoming study that deploys a machine-learning denoiser operating on raw, noise-corrupted traces.

### v2

@@ Abstract
Ultra-high-energy cosmic-ray (UHECR) observatories require unbiased direction reconstruction to enable multi-messenger astronomy with sparse, nanosecond-scale radio pulses. Explicit likelihood methods often rely on simplified models, which may bias results and understate uncertainties. We introduce a simulation-based inference pipeline that couples a physics-informed graph neural network (GNN) to a normalizing-flow posterior within the Learning the Universe Implicit Likelihood Inference framework. Each event is seeded by an analytic plane-wavefront fit; the GNN refines this estimate by learning spatiotemporal correlations among antenna signals, and its frozen embedding conditions an eight-block autoregressive flow that returns the full Bayesian posterior. Trained on about $8,000$ realistic UHECR air-shower simulations generated with the ZHAireS code, the posteriors are temperature-calibrated to meet empirical coverage targets. We demonstrate a sub-degree median angular resolution on test UHECR events, and find that the nominal 68% highest-posterior-density contours capture $71\%\pm 2\%$ of true arrival directions, indicating a mildly conservative uncertainty calibration. This approach provides physically interpretable reconstructions, well-calibrated uncertainties, and rapid inference, making it ideally suited for upcoming experiments targeting highly inclined events, such as GRAND, AugerPrime Radio, and BEACON.
@@ I Introduction
Ultra-high-energy cosmic rays (UHECRs) are the most energetic particles detected in nature Bird and others [1995], Abbasi and others [2023], but their origins and propagation mechanisms remain poorly understood Anchordoqui [2019], Ahlers and Halzen [2015], Alves Batista [2024]. Accurate detection and reconstruction of these rare events have significant implications for fundamental physics and astronomy Alves Batista and others [2019], Bustamante [2024], Guépin et al. [2022]. In particular, accurate arrival directions underpin inference for UHECRs (they set the shower-axis direction and, together with the core position, the event geometry used by downstream reconstructions), so mis-pointing biases the depth of shower maximum and the energy scale that feed composition studies Decoene et al. [2021], Apel and others [2014], Schlüter et al. [2020]. For ultra-high-energy neutrinos, precise pointing is essential for source association and for identifying Earth-skimming trajectories that arrive from below the horizon Schröder [2017].
@@
Next-generation observatories are rising to this challenge by focusing on radio detection techniques, which offer all-sky coverage and high-duty cycles Huege [2016]. Large-scale projects such as the planned Giant Radio Array for Neutrino Detection (GRAND) Álvarez-Muñiz and others [2020], AugerPrime Radio Aab and others [2016], and BEACON Southall and others [2023] are designed to instrument vast areas with antenna networks to capture the nanosecond radio pulses from extensive air showers. These instruments are particularly sensitive to inclined (nearly horizontal) air showers, which produce broad radio footprints on the ground and are a prime channel for detecting Earth-skimming neutrinos and high-energy cosmic rays near the horizon. With their large apertures, radio arrays open a new window for ultra-high-energy neutrino detection. For UHECRs, radio provides high-duty-cycle measurements and can constrain shower development relevant to composition in non- to moderately-inclined geometries. Extending robust composition sensitivity to very inclined showers remains an active area of development and may require hybrid measurements.
@@
The detection of UHECRs using radio arrays is now a well-established technique. Modern radio arrays—such as LOPES Falcke and others [2005], CODALEMA Ardouin and others [2005], LOFAR van Haarlem and others [2013], AERA Huege [2019], and TREND Charrier and others [2019]; reconstruct the incoming direction of UHECRs by fitting a geometric wavefront model to the measured trigger times at known antenna positions.
@@
A reconstruction pipeline has recently been proposed Decoene et al. [2021] for GRAND that leverages the amplitude distribution of the radio signal in a physics-informed, multi-stage approach. The procedure begins with a plane wavefront (PWF) timing fit to obtain an initial estimate of the arrival direction. This initial estimate is subsequently refined by introducing a mild curvature represented by a spheroidal surface wavefront model. In the final step, the lateral distribution of the radio amplitude is introduced to constrain the shower core position and axis orientation precisely. This hierarchical approach yields robust reconstruction performance, especially for air showers at high zenith angles Guelfand et al. [2025].
@@
Although extremely promising, this method also exposes the limits of current reconstruction pipelines. Its successive steps (i.e., planar fitting, curvature correction, and amplitude weighting) rely on model-dependent assumptions that could introduce biases. Traditional explicit likelihood methods, such as formulating a likelihood for observed arrival times given a trial geometry and exploring it via Markov Chain Monte Carlo, are conceptually appealing but practically prohibitive for large arrays and complex wavefront structures Guelfand et al. [2025]. Template‐matching schemes confront similar scalability bottlenecks; reproducing the full span of arrival directions, primary energies, and atmospheric states demands vast libraries of simulated air-shower footprints, whose storage and compute requirements grow steeply with each gain in angular or temporal resolution [Desmet et al., 2025].
@@
Equally important, conventional reconstruction algorithms often fail to rigorously quantify uncertainties; while they may yield a best-fit arrival direction, recovering the full posterior distribution remains non-trivial and is frequently overlooked. The recent analytical PWF solution in Ref. Ferrière et al. [2025] addresses part of this gap by providing a closed-form direction estimator with error estimates for the planar approximation. Yet, a planar wavefront model can only capture so much. To fully exploit the rich information in the radio signal pattern, more flexible inference approaches are needed.
@@
A promising route to overcome the limitations of explicit likelihood methods is simulation-based inference (SBI) Cranmer et al. [2020], also known as implicit or likelihood-free inference. Rather than relying on an analytic likelihood function, SBI leverages high-fidelity simulations of the physics process to directly learn the relationship between model parameters (e.g. shower direction) and observables (antenna times and signals). In a Bayesian context, SBI methods aim to directly estimate the posterior distribution of parameters by training machine learning models on simulated data, thereby “learning” an approximate likelihood function or posterior emulator (see Ref. Miller et al. [2020], Ho and others [2024] and references therein).
This approach is well suited to astroparticle problems like UHECR air showers, where trigger times at fixed stations can be modeled with simple parametric likelihoods under geometric wavefront models (e.g., plane or hyperbolic with Gaussian jitter) Apel and others [2014], van Haarlem and others [2013], Huege [2019]. However, a tractable closed-form likelihood for the full radio data vector (multi-polarization waveforms across many stations, non-Gaussian and correlated noise, thresholds, and array-response systematics) is lacking. High-fidelity forward models such as CoREAS Huege [2016] and ZHAireS Alvarez-Muñiz et al. [2012] capture these effects only through simulation. Simulation-based inference leverages such simulators to perform Bayesian inference without prescribing an explicit analytic likelihood, thereby avoiding restrictive residual models while keeping assumptions transparent in the simulator and prior Cranmer et al. [2020].
@@
In recent years, SBI techniques have gained traction across astrophysics and cosmology, often outperforming traditional analysis methods Mishra-Sharma and Cranmer [2022], Montel et al. [2022], Tucci and Schmidt [2024], Modi et al. [2024], Franco Abellán et al. [2024], Bhardwaj et al. [2023], Karchev and Trotta [2024], Savchenko et al. [2025], Eckner et al. [2025], List et al. [2025]. Examples range from inferring cosmological parameters from galaxy surveys to decoding gravitational wave signals, where neural density estimators have achieved faster and more accurate results than classical MCMC in high-dimensional parameter spaces (See also Ref. Miller et al. [2020] for a detailed discussion). These successes underscore the appeal of SBI; it provides fully Bayesian posteriors with uncertainty quantification for problems that are otherwise intractable with explicit likelihoods Ho and others [2024]. Importantly, likelihood-free approaches can naturally handle complicated selection effects or “cuts” on the data by applying the same cuts to simulated data, a notable advantage when dealing with complex experimental setups.
@@
One such framework is Learning the Universe—Implicit Likelihood Inference (LtU-ILI) Ho and others [2024]. LtU-ILI provides an end-to-end SBI platform that incorporates adaptable neural encoders, inference strategies, and rigorous validation tools. A notable feature of LtU-ILI is its support for complex data representations through custom “embedding” networks—including convolutional networks for image-like data and graph neural networks (GNNs) for relational data.
@@
This is particularly pertinent for radio detector arrays such as GRAND, where the spatial configuration of antennas and their trigger times can be naturally modeled as a graph (with antennas as nodes and edges connecting neighbor detectors). Such a representation enables graph neural networks to effectively capture spatiotemporal correlations in the developing air-shower front. Moreover, embedded in a Bayesian workflow, LtU-ILI audits its posteriors with coverage tests to verify statistical calibration [Ho and others, 2024]. The result is an inference pipeline that remains computationally tractable while preserving clear physical meaning.
@@ II.1 Physical Modeling of Extensive Air Showers
When an UHECR primary strikes the atmosphere, it initiates a cascade of secondary particles known as an extensive air shower (EAS). The primary collision produces a multitude of daughter particles that branch into a hadronic component (secondary mesons like pions and their decay products) and an electromagnetic component (e.g., photons from $\pi^{0}$ decay feeding $e^{+}e^{-}$ cascades) Huege [2016]. This process continues until the primary’s energy is dissipated, yielding a shower of billions of particles spread over kilometers by the time it reaches ground level. Accurately modeling these cascades is essential, as the spatial, temporal, and energy distributions of the secondaries encode the nature of the primary particle Alvarez-Muñiz et al. [2012]. In particular, the arrival direction of the primary can be inferred from the pattern of particle arrival times or atmospheric emissions, providing critical clues to the UHECR’s origin Decoene et al. [2021], Guelfand et al. [2025], Ferrière et al. [2025].
Our work focuses on radio-detection of air showers, a technique that has matured over the past two decades. As the EAS evolves, its charged component experiences geomagnetic deflection and charge separation in the Earth’s magnetic field, driving transverse currents that emit coherent radio pulses. Additionally, a negative charge excess builds up in the shower front (due to electrons knocking out ambient electrons from air molecules), leading to Askaryan radiation Huege [2016], Schröder [2017]. The total radio signal is a superposition of these geomagnetic and Askaryan contributions. Modern experiments like LOFAR van Haarlem and others [2013], AERA Huege [2019], and others have demonstrated that radio arrays can measure UHECR showers with impressive precision, rivaling conventional particle detectors in reconstructing the shower’s arrival direction and even its energy and shower maximum ($X_{\rm max}$). Radio detection thus provides a robust method to observe UHECR events over extensive areas and under most weather conditions (with the important exception of thunderstorms periods). This is achieved by recording the nanosecond-scale electromagnetic pulses generated by air showers using broadband antennas such as log–periodic dipole array, butterfly/fat-dipole designs, and inverted-V or crossed-dipole elements Aab and others [2016], Huege [2019], van Haarlem and others [2013], Falcke and others [2005], Charrier and others [2019], Álvarez-Muñiz and others [2020], Southall and others [2023].
@@ II.2 Synthetic Data Generation and Radio-Signal Formation
To infer the arrival directions of UHECRs, we utilize detailed Monte Carlo simulations of extensive air showers. These simulations Macias et al. [2024] were developed by the GRAND collaboration to facilitate the design and optimization of detection algorithms for GRANDproto300 Chiche [2024], a prototype array comprising $\sim 300$ antennas, presently under construction in China’s Gobi Desert.
@@
The simulations were constructed with the ZHAireS code  [^1] http://aires.fisica.unlp.edu.ar/zhaires/, a simulation package built on the AIRES Halzen et al. [1991], Zas et al. [1992] air-shower engine and extended to calculate radio emission in detail. ZHAireS Alvarez-Muñiz et al. [2012] tracks the full shower development in the atmosphere and computes the radiated electromagnetic fields from charged particle motion based on first principles. Each simulation propagates particles through the atmosphere in small steps, using Liénard-Wiechert potentials to calculate the emitted electric field from each track segment.
@@
This simulator inherently accounts for interference effects among the myriad moving charges and includes realistic propagation conditions, such as the altitude-dependent atmospheric refractive index. As a result, ZHAireS can predict the complete radio pulse waveform at any observer position, in both time and frequency domains. In addition, ZHAireS can simulate the response of a given antenna array to the shower: for each simulated event, it outputs the electric field trace expected at each antenna, from which one can extract observables like signal peak times and electric field (E-field) amplitudes (see, e.g., Ref. Sanchez et al. [2025] and references therein).
@@
We generated time-domain electric-field traces at ground observer points with ZHAireS for downward-going air showers initiated by protons and iron nuclei, spanning energies from 0.4 to 4 EeV; no antenna/electronics response is applied. Our simulations span zenith angles $37^{\circ}\leq\theta\leq 87^{\circ}$ with uniform coverage in azimuth ($0^{\circ}$$360^{\circ}$) and adopt the GRANDProto300 geomagnetic field and atmospheric model.
To reproduce the timing accuracy of the GRANDproto300 digitizer, we take antenna trigger times directly from the clean ZHAireS electric-field traces (no frequency bandpass applied). For each polarization component, we apply a Hilbert transform and use its magnitude to define the envelope. The trigger time for each antenna is taken as the instant when the Hilbert envelope reaches its maximum, using the polarization channel with the largest peak as reference. Antennas whose envelope amplitude falls below $60~\mu\mathrm{V/m}$ are discarded, and only air showers detected by at least five antennas are retained. We add a Gaussian jitter of width $\sigma=5~\mathrm{ns}$ to each trigger time (matching the GPS synchronization uncertainty measured for GRANDproto300 Alves Batista and others [2025]). This yields a clean set of geometric timing measurements while keeping amplitude effects separate (the network is trained only on station coordinates and arrival times x,y,z,t{x,y,z,t}).
Because we do not apply a hardware bandpass here, this $60~\mu\mathrm{V/m}$ envelope cut should not be interpreted as a realistic trigger threshold tied to a specific 50-200 MHz noise floor; it is used only as a simulation-level quality cut to define which antennas provide timing input. A full treatment of amplitude and noise contributions (including bandpass response and gain calibration) will be addressed in a dedicated amplitude-aware study in follow-up work.